OpenStudy (anonymous):

Let the random variable X be the outcome of rolling a 6-sided die. Assuming the die is fair, the probability distribution is as follows: X 1 2 3 4 5 6 Probability 1/6 1/6 1/6 1/6 1/6 1/6 Using the above table,the probability that X is less than or equal to 2 is: A.1/6 B.1/5 C.1/3 D.1/2 E.2/3

OpenStudy (anonymous):

is it C? i just need confirmation @perl

OpenStudy (perl):

can you show me your reasoning

OpenStudy (anonymous):

1 and 2 are the only ones less than or equal to 2 so it would be 2/6 and simplifies to 1/3

OpenStudy (perl):

P( X<=2 ) = P( X =1 or X=2) = P(X=1) + P(X=2) = 1/6 + 1/6

OpenStudy (perl):

so you did it right .

OpenStudy (anonymous):

okay great thank you! i need help with another is that okay?

OpenStudy (perl):

ok

OpenStudy (anonymous):

Which of the following represents the sample space for rolling a 6-sided die and tossing a coin? A. {1H, 2H, 3H, 4H, 5H, 6H, 1T, 2T, 3T, 4T, 5T, 6T} B. {1, 2, 3, 4, 5, 6} C. {H, T} D. {1, 2, 3, 4, 5, 6, H, T}

OpenStudy (perl):

what is your intuition say

OpenStudy (anonymous):

i think its B because a sample space is the set of all possible outcomes

OpenStudy (anonymous):

*D

OpenStudy (perl):

does b) mention a coin, heads or tails?

OpenStudy (anonymous):

i meant D

OpenStudy (perl):

almost. we want to show you roll a die *and* a coin. so for example 1H would be a sample point.

OpenStudy (anonymous):

so it would A?

OpenStudy (perl):

you can think of them as simultaneous events, you roll a die at the same time as flipping a coin

OpenStudy (perl):

right

OpenStudy (anonymous):

okay thank you! can you check my work on this? If P (A) = 0.2 and P (B) = 0.3 and A and B are independent but NOT disjoint, find P (A or B). A 0.06 B 0.10 C 0.44 D 0.50 E Cannot be determined from the information given. i say A because its independent you do (.2)(.3) = 0.06

OpenStudy (anonymous):

what do you think? i say A because its independent you do (.2)(.3) = 0.06

OpenStudy (anonymous):

@perl

OpenStudy (perl):

one moment

OpenStudy (anonymous):

ok

OpenStudy (perl):

The general 'or' formula is : P(A or B ) = P(A) + P(B) - P( A & B )

OpenStudy (anonymous):

i get 0.44

OpenStudy (perl):

okay thank you! can you check my work on this? If P (A) = 0.2 and P (B) = 0.3 and A and B are independent but NOT disjoint, find P (A or B). P(A or B) = P(A) + P(B) - P(A & B ) because A,B are independent P(A or B) = P(A) + P(B) - P(A) *P(B) = 0.2 + 0.3 - (0.2)*(0.3)

OpenStudy (perl):

yes thats correct

OpenStudy (anonymous):

okay sorry can you check one more?

OpenStudy (anonymous):

Let the random variable X be the outcome of rolling a 6-sided die. Assuming the die is fair, the probability distribution is as follows: X 1 2 3 4 5 6 Probability 1/6 1/6 1/6 1/6 1/6 1/6 Using the table of probabilities above,what is the probability of rolling a value greater than 3? A. 1/6 B. 1/4 C. 1/3 D. 1/2 E. 1

OpenStudy (anonymous):

I said 1/6 because the probability of getting any number 1/6 or would it be 1/3 because we leave the numbers 3 and under out

OpenStudy (perl):

P(X > 3) = P( X = 4 or X = 5 or X = 6) = P( X = 4) + P(X = 5) + P( X = 6) = 1/6 + 1/6 + 1/6

OpenStudy (anonymous):

so 1/3 is correct?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

yes

OpenStudy (anonymous):

okay thank you so much!! you're amazing as always

OpenStudy (perl):

Your welcome .